# Colorful Pinball: Density-Weighted Quantile Regression for Conditional Guarantee of Conformal Prediction

### v3

@@ 4.1 Algorithm details
$\phantom{=}\mathbb{E}_{X}[f_{S\mid X}(q_{\tau}(X))\mathcal{E}(X)]=\mathbb{E}_{X,Y}\left[f_{S\mid X}(q_{\tau}(X))\rho_{\tau}(\hat{q}_{\tau}(X),S(X,Y))\right]+\mathrm{const}$ (17)
@@
6: Fine-tune the parameters of $\hat{q}_{\text{main}}$ on $\mathcal{D}_{\text{cal},2}$ to get the final quantile estimator $\hat{q}_{\tau}$.
@@ 4.2 Towards better finite-sample stability
  • Loss Mixing: We modify the optimization objective as a convex combination of the normalized weighted pinball loss and the plain pinball loss.
In essence, these two strategies correspond to imposing artificial upper and lower bounds on the estimated weights, a modification motivated by the theoretical bottlenecks identified in the proof of our main theorem (Theorem 5.2). Furthermore, such clipping strategies are widely recognized as effective in settings involving reciprocal weighting; they significantly reduce variance at the cost of introducing slight bias, a trade-off commonly utilized in policy evaluation (e.g., clipping inverse propensity weights; Swaminathan and Joachims, 2015). We demonstrate the empirical benefits of these strategies in Section 6 and Appendix A.
@@ 5 Theoretical Analysis
To approximate MSCE, i.e., the MSE of coverage probability $\pi(x)$, we employ a two-step transformation. First, we substitute the target $\pi(x)$ with $F_{S\mid X}(\hat{q}_{\tau}(x))$; second, we select the leading term derived in Proposition 3.4 as our final optimization surrogate. According to Algorithm 1, the first substitution introduces a discrepancy due to the global conformal shift $\hat{\gamma}$ required for marginal coverage. However, as shown by Plassier et al. [2025a], this gap decays exponentially with the calibration sample size mm [^1] We denote $|\mathcal{D}_{\text{calib},3}|=m$ and $|\mathcal{D}_{\text{calib},1}|=|\mathcal{D}_{\text{calib},2}|=n$., provided $\hat{q}_{\tau}$ satisfies certain regularity conditions that controls the quality of quantile estimator $\hat{q}_{\tau}$ (see Appendix C.1). Intuitively, with a moderately large sample size of $\mathcal{D}_{\text{cal}}$ (e.g., in the hundreds), we can focus on characterizing the finite-sample performance of our method through analyzing the MSE of $F_{S\mid X}\left(\hat{q}_{\tau}(x)\right)$, and naturally the surrogate, expected risk of density-weighted pinball loss.
@@
    Local Hölder-type norm equivalence. There exist constants $C_{\mathrm{norm}}>0$, $\nu\in(0,1]$, and $r>0$ such that for all $g\in\mathcal{G}$ satisfying $\|g-g^{\star}\|_{L_{2}(\mathbb{P}_{X})}\leq r$,
@@
Among these assumptions, the first two are standard technical conditions, whereas the third is less conventional and warrants a more detailed discussion due to its pivotal role in our proof. The primary challenge in bounding the generalization error of our method stems from the estimated reciprocal weights: specifically, we require a pointwise lower bound on the denominator, whereas typical guarantees yield only $L_{2}$ convergence. To bridge this gap, we introduce this norm-equivalence assumption, which we relax to hold only locally within a neighborhood of the true quantile function $g^{*}$. In general, it suffices that $g-g^{*}$ belongs to a Hölder class for this assumption to hold. We provide a proof of this claim in Appendix B.4, which can also be obtained as a special case of the celebrated Gagliardo–Nirenberg interpolation inequality. Finally, we note that the exponent $\nu$ increases with the smoothness of the function class; for instance, $\nu=1$ for parametric models or regression in a Reproducing Kernel Hilbert Space with a smooth kernel (e.g., Gaussian or Laplacian).
@@
Under Assumption 5.1, with probability at least $1-3\zeta$, setting the finite-difference bandwidth to $\delta^{\star}\asymp(\mathfrak{R}_{n}(\mathcal{G}))^{1/3}$, with appropriately large $n$, the excess risk satisfies:
@@
$\mathcal{R}(\hat{g})-\mathcal{R}\left(g^{\star}\right)$ $\leq C_{1}\mathfrak{R}_{n}(\mathcal{G})+C_{2}\sqrt{\frac{\log(1/\zeta)}{n}}$ (26)
$+C_{3}\left(\mathcal{E}_{q}(n)^{2}+C_{4}\mathcal{E}_{q}(n)^{1+\nu}\sqrt{\frac{\log(1/\zeta)}{n}}\right$.
@@
where constants $C_{1}=4B_{w}L_{\rho}$, $C_{2}=2\sqrt{2}B_{w}L_{\rho}\sigma_{S}$, $C_{3}=2^{5/3}3^{2/3}M_{\rho,\mathcal{G}}B_{w}^{2}(B_{q}^{\prime\prime\prime})^{1/3}$, $C_{4}=2^{\nu+1/2}C_{\mathrm{norm}}$, $C_{5}=C_{\mathrm{norm}}^{2}/3$, $L_{\rho}=\max\{\tau,1-\tau\}$.
In the asymptotic regime, there is a rate loss due to weight estimation, compared to the direct quantile regression for $q_{\tau}$ with plain pinball loss. Compared to direct quantile regression with the pinball loss, our method exhibits a slower asymptotic rate due to the additional error introduced by estimating density-based weights. However, we emphasize that the primary value of conformal prediction lies in the finite-sample regime, which motivates the exact non-asymptotic expression in our theory. In finite samples, although direct quantile regression may converge faster, the misalignment of its optimization objective with conditional coverage can substantially limit its performance, as illustrated in Proposition 3.4 and Example 3.1. On the other hand, when higher-order central finite-difference schemes of order $k$ are employed Fornberg [1988], the resulting excess risk rate can be further improved to $O(\mathfrak{R}_{n}(\mathcal{G})^{2k/(2k+1)})$.
6 Experiments
Data.
We consider eight real-world regression benchmarks spanning engineering systems, energy, transportation, and economics. Among them, three datasets (Bike, Diamond, and Superconductivity) have one-dimensional responses, while the remaining five involve multi-dimensional responses. The covariate dimension ranges from a moderate scale (tens of features), and the sample size varies between 10k and 50k. Each dataset is randomly split into training, calibration, and test sets with a ratio of 6:2:2.
Baselines.
We consider standard baselines including Split Conformal Prediction (Split; Papadopoulos et al., 2002), Partition Learning Conformal Prediction (PLCP; Kiyani et al., 2024), Gaussian Scoring Braun et al. [2025a], CQR Romano et al. [2019], RCP Plassier et al. [2025a]. In addition, to separately validate the value of our approximation described in Proposition 3.4, we select a parametric likelihood on nonconformity score, the Asymmetric Laplace Distribution (ALD), of which the maximum likelihood estimation (MLE) is equivalent to minimizing scaled pinball loss and a regularization term on scales $\sigma(x)$. We construct two baselines based on CQR and RCP through substituting the quantile regression with MLE of ALD. We provide the details of baselines in Appendix A.2.
Except for Gaussian Scoring, we adopt $l_{\infty}$ norm of residuals as nonconformity scores Diquigiovanni et al. [2024], which reduces to the standard absolute residual score for the case of one-dimensional labels.
Metrics.
To evaluate the conditional coverage in real-world datasets, where we have no access to the data generation mechanism, we adopt MSCE (approximated by K-means), Worst-Slice Coverage (WSC; Cauchois et al., 2021, Romano et al., 2020), excess risk of the target coverage ($\ell$-ERT; Braun et al., 2025b). In general, given marginal validity, better concentration of conditional coverage probabilities is indicated by lower MSCE, higher WSC, and lower $\ell$-ERT. Moreover, WSC also serves as a robustness measure, characterizing the worst-case coverage. Throughout the experiments, we set the target coverage level $\tau=90\%$. We also report the mean (per dimension) of the log volume of the predictive set to characterize the efficiency. All methods are run with 20 Monte Carlo repetitions for evaluation, and we report $\operatorname{mean}\pm\operatorname{std}$. Metric details and related results are provided in Appendices A.4 and A.5, respectively.
Table 1: Comparison of Mean Squared Coverage Error (MSCE, $\downarrow$). Best results are bolded.
@@ 6 Experiments
Table 2: Comparison of Worst-Slice Coverage (WSC, $\uparrow$). Best results are bolded.
@@
Table 3: Comparison of $\ell_{1}$ Excess Risk of the Target coverage ($\ell_{1}$-ERT, $\downarrow$). Best results are bolded.
@@ 6.1 Results Analysis
Regarding the ablation studies, we first observe that the parametric likelihood (ALD) yields consistent benefits, further validating the approximation in Proposition 3.4. Second, our two additional stability mechanisms demonstrate high effectiveness. Finally, as shown in Appendix A.5, the influence of $\delta$ is relatively minor (comparing our default $\delta=2\%$ against $1\%$ and $5\%$). Furthermore, the limited benefit observed from pure multitasking (co-training three quantiles) further highlights the significance of our specific methodological contributions.
@@ A.1 Dataset description
We introduce our benchmarks in alphabetical order, following the sequence presented in our tables. First, we provide a summary of all eight datasets, after which we detail each one individually. We note that the majority of these datasets are sourced from the UCI repository Kelly et al. [2025].
@@ A.2 Baseline details
$h^{*},\boldsymbol{q}^{*}=\underset{\begin{subarray}{c}\boldsymbol{q}\in\mathbb{R}^{G},h\in\mathcal{H}\end{subarray}}{\operatorname{argmin}}\frac{1}{n}\sum_{j=1}^{n}\sum_{i=1}^{G}h^{i}\left(X_{j}\right)\rho_{\tau}\left(q_{i},S_{j}\right),$ (28)
@@
As described in Braun et al. [2025a], this method assumes a multivariate Gaussian density for $Y\mid X$. On the training set, we train two MLPs to simultaneously predict the conditional mean $\hat{\mu}(X)$ and covariance matrix $\hat{\Sigma}(X)$ by minimizing the negative log-likelihood (NLL). The nonconformity score is defined as the Mahalanobis distance:
@@
which reduces to the standard normalized residual $S(X,Y)=|Y-\hat{\mu}(X)|/\hat{\sigma}(X)$ Lei et al. [2018] when $Y$ is one-dimensional.
@@ A.3 CPCP details
Although we derive the non-asymptotic expression $\delta(n)$, we fix $\delta=0.02$ throughout our experiments to avoid cherry-picking. In general, a smaller $\delta$ reduces the bias of the finite-difference estimation but requires a larger sample size to maintain stability. Our theory establishes that the optimal rate for this trade-off is $\delta^{\star}\asymp O(n^{-1/6})$.
@@ A.4 Metric details
For MSCE, we use K-means to partition samples into $K$ groups Braun et al. [2025b]. There is a fundamental trade-off here: increasing $K$ approximates the conditional coverage with finer granularity, but reduces the sample size within each group, thereby decreasing the precision of the empirical coverage estimate. Our guiding principle is to ensure sufficient samples (on the order of hundreds) within each group to evaluate coverage with percent-level precision. Consequently, we examine two settings: $K=10$ and $K=30$, as reported in Tables 5 and 6. We note that this approach depends on Euclidean distance, which becomes increasingly unreliable in high dimensions, rendering it ineffective for high-dimensional $X$.
@@
Romano et al. [2020], Cauchois et al. [2021] propose the worst-slab coverage (WSC). We first define a slab in the covariate space as:
@@
In practice, we estimate WSC for a given $\mathcal{C}_{\alpha}$ by sampling 1,000 independent vectors $v$ from the unit sphere in $\mathbb{R}^{d}$ and optimizing parameters $a$ and $b$ via grid search. To mitigate finite-sample negative bias, we partition the test data into two disjoint subsets (e.g., with a $25\%$-$75\%$ split). We use the first subset to determine the optimal parameters $v^{*},a^{*},b^{*}$, and the second to evaluate the conditional coverage:
$\mathbb{P}\left[Y\in\mathcal{C}_{\alpha}(X)\mid X\in S_{v^{*},a^{*},b^{*}}\right].$
@@
Braun et al. [2025b] introduce the Excess Risk of the Target coverage (ERT), which casts coverage evaluation as a binary classification problem. This approach relies on the insight that under perfect conditional coverage, the coverage indicator $Z=\mathbb{I}\{Y\in\mathcal{C}_{\alpha}(X)\}$ behaves as a Bernoulli variable with constant probability $1-\alpha$, independent of $X$. Consequently, any classifier $h(X)$ that achieves a lower risk than the constant predictor $1-\alpha$ exposes a violation of conditional validity. The metric is defined as the risk reduction:
@@
where $\mathcal{R}_{\ell}$ is the risk under a proper loss function $\ell$. Depending on the choice of $\ell$, this metric estimates different distances; specifically, we consider $\ell_{1}$-ERT and $\ell_{2}$-ERT, which provide lower-bound estimates for $\mathbb{E}[|\pi(X)-\tau|]$ and $\mathbb{E}[(\pi(X)-\tau)^{2}]$ (i.e., MSCE), respectively. In our experiments, we employ Logistic Regression as the classifier $h$ to estimate these quantities.

### v4

@@ 4.1 Algorithm Details
$\phantom{=}\mathbb{E}_{X}[f_{S\mid X}(q_{\tau}(X))\mathcal{E}(X)]$ (17)
$=\mathbb{E}_{X,Y}\left[f_{S\mid X}(q_{\tau}(X))\rho_{\tau}(\hat{q}_{\tau}(X),S)\right]+\mathrm{const}$
@@
6: Fine-tune the $\hat{q}_{\text{main}}$ on $\mathcal{D}_{\text{cal},2}$ to get the final quantile estimator $\hat{q}_{\tau}$ (freezing backbone and auxiliary heads).
@@ 4.2 Towards Better Finite-Sample Stability
  • Loss Mixing: We modify the optimization objective as a convex combination of the (self-normalized) weighted pinball loss and the standard pinball loss.
In essence, these two strategies correspond to imposing artificial upper and lower bounds on the estimated weights, a modification motivated by the theoretical bottlenecks identified in the proof of our main theorem (Theorem 5.2). Furthermore, such clipping strategies are widely recognized as effective in settings involving reciprocal weighting; they significantly reduce variance at the cost of introducing slight bias, a trade-off commonly utilized in policy evaluation (e.g., clipping inverse propensity weights; Swaminathan and Joachims, 2015). We demonstrate the empirical benefits of these strategies in Section 6 and Appendix A.5.
@@ 5 Theoretical Analysis
To approximate MSCE, i.e., the MSE of coverage probability $\pi(x)$, we employ a two-step transformation. First, we substitute the target $\pi(x)$ with $F_{S\mid X}(\hat{q}_{\tau}(x))$; second, we select the leading term derived in Proposition 3.4 as our final optimization surrogate. According to Algorithm 1, the first substitution introduces a discrepancy due to the global conformal shift $\hat{\gamma}$ required for marginal coverage. However, as shown by Plassier et al. (2025a), this gap decays exponentially with the calibration sample size mm [^2] We denote $|\mathcal{D}_{\text{calib},3}|=m$ and $|\mathcal{D}_{\text{calib},1}|=|\mathcal{D}_{\text{calib},2}|=n$., provided $\hat{q}_{\tau}$ satisfies certain regularity conditions that controls the quality of quantile estimator $\hat{q}_{\tau}$ (see Appendix C.1). Intuitively, with a moderately large sample size of $\mathcal{D}_{\text{cal}}$ (e.g., in the hundreds), we can focus on characterizing the finite-sample performance of our method through analyzing the MSE of $F_{S\mid X}\left(\hat{q}_{\tau}(x)\right)$, and naturally the surrogate, expected risk of density-weighted pinball loss.
@@
    Local Hölder-type norm equivalence. There exist constants $C_{\mathrm{norm}}>0$, $\nu\in[1/3,1]$, and $r>0$ such that for all $g\in\mathcal{G}$ satisfying $\|g-g^{\star}\|_{L_{2}(\mathbb{P}_{X})}\leq r$,
@@
Among these assumptions, the third is less conventional and warrants a more detailed discussion due to its pivotal role in our proof. The primary challenge in bounding the generalization error of our method stems from the estimated reciprocal weights: specifically, we require a pointwise lower bound on the denominator, whereas typical guarantees yield only $L_{2}$ convergence. To bridge this gap, we introduce this norm-equivalence assumption, which we relax to hold only locally within a neighborhood of the true quantile function $g^{\star}$. In general, it suffices that $g-g^{\star}$ belongs to a Hölder class for this assumption to hold. We provide a proof of this claim in Appendix B.4, which can also be obtained as a special case of the celebrated Gagliardo–Nirenberg interpolation inequality. Finally, we note that the exponent $\nu$ increases with the smoothness of the function class; for instance, $\nu=1$ for parametric models or regression in a Reproducing Kernel Hilbert Space with a smooth kernel.
@@
Under Assumption 5.1, with probability at least $1-3\zeta$, setting the finite-difference bandwidth to $\delta^{\star}\asymp(\mathfrak{R}_{n}(\mathcal{G}))^{1/3}$, with appropriately large $n$, we have:
@@
[…]
$+C_{3}M_{\rho,\mathcal{G}}\left(\mathcal{E}_{q}(n)^{2}+C_{4}\mathcal{E}_{q}(n)^{1+\nu}\sqrt{\frac{\log(1/\zeta)}{n}}\right$.
@@
where constants $C_{1}=4B_{w}L_{\rho}$, $C_{2}=2\sqrt{2}B_{w}L_{\rho}\sigma_{S}$, $C_{3}=4\times 3^{2/3}B_{w}^{2}(B_{q}^{\prime\prime\prime})^{1/3}$, $C_{4}=2^{\nu-1/2}C_{\mathrm{norm}}$, $C_{5}=C_{\mathrm{norm}}^{2}/6$, $L_{\rho}=\max\{\tau,1-\tau\}$.
In the asymptotic regime, there is a rate loss due to weight estimation, compared to the direct quantile regression for $q_{\tau}$ with plain pinball loss. Compared to direct quantile regression with the pinball loss, our method exhibits a slower asymptotic rate due to the additional error introduced by estimating density-based weights. However, we emphasize that the primary value of conformal prediction lies in the finite-sample regime, which motivates the exact non-asymptotic expression in our theory. In finite samples, although direct quantile regression may converge faster, the misalignment of its optimization objective with conditional coverage can substantially limit its performance, as illustrated in Proposition 3.4 and Example 3.1. On the other hand, when higher-order central finite-difference schemes of order $k$ are employed (Fornberg, 1988), the resulting excess risk rate can be further improved to $O(\mathfrak{R}_{n}(\mathcal{G})^{2k/(2k+1)})$.
Table 1: Comparison of MSCE. Mean Squared Conditional Error (MSCE, $\downarrow$) on 8 high-dimensional benchmarks. Best results are bolded.
@@
Table 2: Comparison of WSC. Worst Slab Coverage (WSC, $\uparrow$) on 8 high-dimensional benchmarks. Best results are bolded.
@@
Table 3: Comparison of $L_{1}$-ERT. $L_{1}$ Excess Risk of the Target coverage ($L_{1}$-ERT, $\downarrow$) on 8 benchmarks. Best results are bolded.
@@
6 Experiments
Data. We consider eight real-world regression benchmarks spanning engineering systems, energy, transportation, and economics. Among them, three datasets (Bike, Diamond, and Superconductivity) have one-dimensional responses, while the remaining five involve multi-dimensional responses. The covariate dimension ranges from a moderate scale (tens of features), and the sample size varies between 10k and 70k. Each dataset is randomly split into training, calibration, and test sets with a ratio of 6:2:2.
Baselines. We consider standard baselines including Split Conformal Prediction (Split; Papadopoulos et al., 2002), Partition Learning Conformal Prediction (PLCP; Kiyani et al., 2024), Gaussian Scoring (Braun et al., 2025a), CQR (Romano et al., 2019), RCP (Plassier et al., 2025a). In addition, to separately validate the value of our approximation described in Proposition 3.4, we select a parametric likelihood on nonconformity score, the Asymmetric Laplace Distribution (ALD), of which the maximum likelihood estimation (MLE) is equivalent to minimizing scaled pinball loss and a regularization term on scales $\sigma(x)$. We construct two baselines based on CQR and RCP through substituting the quantile regression with MLE of ALD. We provide the details of baselines in Appendix A.2.
Except for Gaussian Scoring, we adopt $l_{\infty}$ norm of residuals as nonconformity scores (Diquigiovanni et al., 2024), which reduces to the standard absolute residual score for the case of one-dimensional labels.
Metrics. To evaluate the conditional coverage in real-world datasets, where we have no access to the data generation mechanism, we adopt MSCE (approximated by K-means), Worst-Slice Coverage (WSC; Cauchois et al., 2021; Romano et al., 2020), excess risk of the target coverage ($\ell$-ERT; Braun et al., 2025b). In general, given marginal validity, better concentration of conditional coverage probabilities is indicated by lower MSCE, higher WSC, and lower $\ell$-ERT. Moreover, WSC also serves as a robustness measure, characterizing the worst-case coverage. Throughout the experiments, we set the target coverage level $\tau=90\%$. We also report the mean (per dimension) of the log volume of the predictive set to characterize the efficiency. All methods are run with 20 Monte Carlo repetitions for evaluation, and we report $\operatorname{mean}\pm\operatorname{std}$. Metric details and related results are provided in Appendices A.4 and A.5, respectively.
@@ 6.1 Results Analysis
Regarding the ablation studies, we first observe that the parametric likelihood (ALD) yields consistent benefits, further validating the approximation in Proposition 3.4. Second, our two additional stability mechanisms demonstrate high effectiveness. Third, as shown in Appendix A.5, the influence of $\delta$ is relatively minor (comparing our default $\delta=2\%$ against $1\%$ and $5\%$). Furthermore, the limited benefit observed from pure multitasking (co-training three heads) indicates that the performance gains are primarily attributable to our core methodology. This finding also underscores our framework’s flexibility, as the quantile regressor can be substituted with alternative predictive models beyond neural networks.
@@ A.1 Dataset Description
We introduce our benchmarks in alphabetical order, following the sequence presented in our tables. First, we provide a summary of all eight datasets, after which we detail each one individually. We note that the majority of these datasets are sourced from the UCI repository (Kelly et al., 2025).
@@ A.2 Baseline Details
$h^{\star},\boldsymbol{q}^{\star}=\underset{\begin{subarray}{c}\boldsymbol{q}\in\mathbb{R}^{G},h\in\mathcal{H}\end{subarray}}{\operatorname{argmin}}\frac{1}{n}\sum_{j=1}^{n}\sum_{i=1}^{G}h^{i}\left(X_{j}\right)\rho_{\tau}\left(q_{i},S_{j}\right),$ (28)
@@
As described in (Braun et al., 2025a), this method assumes a multivariate Gaussian density for $Y\mid X$. On the training set, we train two MLPs to simultaneously predict the conditional mean $\hat{\mu}(X)$ and covariance matrix $\hat{\Sigma}(X)$ by minimizing the negative log-likelihood (NLL). The nonconformity score is defined as the Mahalanobis distance:
@@
which reduces to the standard normalized residual $S(X,Y)=|Y-\hat{\mu}(X)|/\hat{\sigma}(X)$ (Lei et al., 2018) when $Y$ is one-dimensional.
@@ A.3 CPCP Details
Although we derive the non-asymptotic expression $\delta(n)$, we take $\delta=0.02$ as default throughout our experiments. As comparison, we also provide the results with $\delta=0.01,0.05$, which is named as CPCP-$0.01$ and CPCP-$0.05$, respectively.
In general, a smaller $\delta$ reduces the bias of the finite-difference estimation but requires a larger sample size to maintain stability. Our theory establishes that the optimal rate for this trade-off is $\delta^{\star}\asymp O(n^{-1/6})$.
@@ A.4 Metric Details
For MSCE, we use K-means to partition samples into $K$ groups. There is a fundamental trade-off here: increasing $K$ approximates the conditional coverage with finer granularity, but reduces the sample size within each group, thereby decreasing the precision of the empirical coverage estimate. Our guiding principle is to ensure sufficient samples (on the order of hundreds) within each group to evaluate coverage with percent-level precision. Consequently, we examine two settings: $K=10$ and $K=30$, as reported in Tables 5 and 6.
@@
Romano et al. (2020); Cauchois et al. (2021) propose the worst-slab coverage (WSC). We first define a slab in the covariate space as:
@@
In practice, we estimate WSC for a given $\mathcal{C}_{\alpha}$ by sampling 1,000 independent vectors $v$ from the unit sphere in $\mathbb{R}^{d}$ and optimizing parameters $a$ and $b$ via grid search. To mitigate finite-sample negative bias, we partition the test data into two disjoint subsets (e.g., with a $25\%$-$75\%$ split). We use the first subset to determine the optimal parameters $v^{\star},a^{\star},b^{\star}$, and the second to evaluate the conditional coverage:
$\mathbb{P}\left[Y\in\mathcal{C}_{\alpha}(X)\mid X\in S_{v^{\star},a^{\star},b^{\star}}\right].$
@@
Braun et al. (2025b) introduce the Excess Risk of the Target coverage (ERT), which casts coverage evaluation as a binary classification problem. This approach relies on the insight that under perfect conditional coverage, the coverage indicator $Z=\mathbb{I}\{Y\in\mathcal{C}_{\alpha}(X)\}$ behaves as a Bernoulli variable with constant probability $1-\alpha$, independent of $X$. Consequently, any classifier $h(X)$ that achieves a lower risk than the constant predictor $1-\alpha$ exposes a violation of conditional validity. The metric is defined as the risk reduction:
@@
where $\mathcal{R}_{\ell}$ is the risk under a proper loss function $\ell$. Depending on the choice of $\ell$, this metric estimates different distances; specifically, we consider $L_{1}$-ERT and $L_{2}$-ERT, which provide lower-bound estimates for $\mathbb{E}[|\pi(X)-\tau|]$ and $\mathbb{E}[(\pi(X)-\tau)^{2}]$ (i.e., MSCE), respectively. In our experiments, we employ Logistic Regression as the classifier $h$ to estimate these quantities.